# Statistics-encoded tensor network approach in disordered quantum many-body spin chains

### v1

@@ Abstract
Summary and outlook.—We propose SeTN, a tensor network approach for studying chaotic dynamics in general disordered quantum systems. By encoding disorder into an auxiliary layer and performing the disorder averaging separately, SeTN restores spatial translational invariance and enables accurate simulations directly in the thermodynamic limit. From the singular value spectrum of SeTN, we derive and numerically confirm a criterion for efficient disorder encoding, $n\gg\alpha^{2}t^{2}$, which indicates that encoding is more efficient under weak disorder, a regime that is typically chaotic and challenging for conventional TN methods. We compute the leading eigenvalue of the SeTN-derived transfer matrix for the SFF in the disordered TFIM, and show that it accurately captures the scaling behavior of the SFF within the accessible time range. Based on the behavior within this range, we conjecture how the single-eigenvalue dominance crosses over to random-matrix-theory behavior, a key aspect of many-body quantum chaos that deserves further investigation.
@@
This SeTN approach can be readily extended to other disordered models and physical observables that can be represented by multilayer TN representations. Examples include Rényi entropies [chanSolutionMinimalModel2018c, zhouEmergentStatisticalMechanics2019], out-of-time-ordered correlators (OTOCs)[nahumOperatorSpreadingRandom2018c, vonkeyserlingkOperatorHydrodynamicsOTOCs2018a, chanSolutionMinimalModel2018c], and spin-glass order parameters[huseLocalizationprotectedQuantumOrder2013a, pekkerHilbertGlassTransitionNew2014a]. The simple and general structure of SeTN is reminiscent of an effective trajectory pairing mechanism in the Feynman-trajectory picture [vonkeyserlingkOperatorHydrodynamicsOTOCs2018a, zhouEmergentStatisticalMechanics2019, zhouEntanglementMembraneChaotic2020], where disorder averaging cancels most quantum trajectories through destructive interference. SeTN may also serve as a bridge between local averaging schemes and more advanced concepts, such as influence functionals [nguyenCorrelationFunctionsTensor2024, sonnerInfluenceFunctionalManybody2021a, yeConstructingTensorNetwork2021, leroseInfluenceMatrixApproach2021b] and entanglement barriers [rathEntanglementBarrierIts2023], and offers a powerful platform for investigating the impact of disorder on dynamical phenomena such as diffusion and hydrodynamics [khemaniOperatorSpreadingEmergence2018c, rakovszkyDiffusiveHydrodynamicsOutofTimeOrdered2018c]. Overall, SeTN provides a unifying framework for exploring the complex interplay between locality, unitarity, disorder, and chaos in time-independent many-body quantum systems.

### v2

@@ V Summary and outlook
We propose SeTN, a tensor network approach for studying chaotic dynamics in general disordered quantum systems. By encoding disorder into an auxiliary layer and performing the disorder averaging separately, SeTN restores spatial translational invariance and enables accurate simulations directly in the thermodynamic limit through a compact MPO transfer matrix.From the singular value spectrum, we derive and numerically confirm a universal criterion for efficient disorder encoding, $n\gg\alpha^{2}t^{2}$, which indicates that encoding is more efficient under weak disorder, a regime that is typically chaotic and challenging for conventional TN methods, making disorder-averaged transfer matrices numerically feasible for generic Hamiltonians.
We compute the leading eigenvalue of the SeTN-derived transfer matrix for the SFF in the disordered TFIM, and show that it accurately captures the scaling behavior of the SFF within the accessible time range. Based on the behavior observed in this regime, we conjecture that the crossover from single-eigenvalue dominance to random-matrix-theory behavior is associated with an increasing near-degeneracy among the leading transfer-matrix eigenvalues. Clarifying how this near-degeneracy develops and how it controls the emergence and system-size dependence of the Thouless time remains an important open problem for future work.
@@
SeTN is broadly applicable to observables with multilayer TN representations, including Rényi entropies [chanSolutionMinimalModel2018c, zhouEmergentStatisticalMechanics2019], out-of-time-ordered correlators (OTOCs) [nahumOperatorSpreadingRandom2018c, vonkeyserlingkOperatorHydrodynamicsOTOCs2018a, chanSolutionMinimalModel2018c], and spin-glass order parameters [huseLocalizationprotectedQuantumOrder2013a, pekkerHilbertGlassTransitionNew2014a], offering a powerful platform for investigating the impact of disorder on dynamical phenomena such as diffusion and hydrodynamics [khemaniOperatorSpreadingEmergence2018c, rakovszkyDiffusiveHydrodynamicsOutofTimeOrdered2018c]. SeTN also serves as a bridge between local averaging schemes and other concepts, such as Feynman-trajectory picture [vonkeyserlingkOperatorHydrodynamicsOTOCs2018a, zhouEmergentStatisticalMechanics2019, zhouEntanglementMembraneChaotic2020], influence functionals [nguyenCorrelationFunctionsTensor2024, sonnerInfluenceFunctionalManybody2021a, yeConstructingTensorNetwork2021, leroseInfluenceMatrixApproach2021b] and entanglement barriers [rathEntanglementBarrierIts2023]. Overall, our construction provides a unified framework for exploring the interplay of locality, unitarity, disorder, and chaos in static many-body systems. In this sense, SeTN complements recent developments such as emergent symmetries in disorder-averaged dynamics [erpeldingExploitingEmergentSymmetries2025a] and translationally invariant MPO descriptions of thermal ensembles [vervoortExtractingAverageProperties2025], and opens multiple avenues for future investigation.